9.29 Calculate the molar concentration of H3O+ in water solutions with the following

OH- molar concentrations:

a) 1.0 x 10^-5

b) 3.4 x 10^-4

c) 8.3 x 10^-9

d) 0.072

e) 3.2
Answer #1

a)

we have below equation to be used:

[H3O+] = Kw/[OH-]

Kw is dissociation constant of water whose value is 1.0*10^-14 at 25 oC

[H3O+] = (1.0*10^-14)/[OH-]

[H3O+] = (1.0*10^-14)/1.0E-5

[H3O+] = 1.0*10^-9 M

Answer: 1.0*10^-9 M

b)

we have below equation to be used:

[H3O+] = Kw/[OH-]

Kw is dissociation constant of water whose value is 1.0*10^-14 at 25 oC

[H3O+] = (1.0*10^-14)/[OH-]

[H3O+] = (1.0*10^-14)/3.4E-4

[H3O+] = 2.941*10^-11 M

Answer: 2.9*10^-11 M

c)

we have below equation to be used:

[H3O+] = Kw/[OH-]

Kw is dissociation constant of water whose value is 1.0*10^-14 at 25 oC

[H3O+] = (1.0*10^-14)/[OH-]

[H3O+] = (1.0*10^-14)/8.3E-9

[H3O+] = 1.205*10^-6 M

Answer: 1.2*10^-6 M

d)

we have below equation to be used:

[H3O+] = Kw/[OH-]

Kw is dissociation constant of water whose value is 1.0*10^-14 at 25 oC

[H3O+] = (1.0*10^-14)/[OH-]

[H3O+] = (1.0*10^-14)/0.072

[H3O+] = 1.4*10^-13 M

Answer: 1.4*10^-13 M

e)

we have below equation to be used:

[H3O+] = Kw/[OH-]

Kw is dissociation constant of water whose value is 1.0*10^-14 at 25 oC

[H3O+] = (1.0*10^-14)/[OH-]

[H3O+] = (1.0*10^-14)/3.2

[H3O+] = 3.1*10^-15 M

Answer: 3.1*10^-15 M